OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 12m the square root of 21 times m and width the square root of 16 times m, which of the following is the area of the rectangle?

OpenStudy (anonymous):

from where u have taken this question

OpenStudy (anonymous):

i think it is a multiple choice question but the answer is : L = 12sqrt(21m) W = 16m A = L * W = 12sqrt(21m)(16m)=192sqrt(21m^2)

OpenStudy (anonymous):

Are these the sides: \[12m \sqrt{21m}\]and\[\sqrt{16m}\]

OpenStudy (anonymous):

\[\text{Area}=\left(12m \sqrt{21 m}\right) \sqrt{16m} =48 \sqrt{21} m^2 \]

OpenStudy (anonymous):

Thank you for the medal.